B) A 1300 -kg car is pushing an out-of-gear 2200 -kgtruck that has a dead battery. When the driver steps on the accelerator, the drive wheels of the car push horizontally against the ground with a force of 4500 N . The rolling friction of the car can be neglected, but the heavier truck has a rolling friction of 760 N , including the "friction" of turning the truck's drivetrain. What is the acceleration aT of the truck?

C) An 89.0 kg spacewalking astronaut pushes off a 620 kg satellite, exerting a 85.0 N force for the 0.540 s it takes him to straighten his arms. How far apart are the astronaut and the satellite after 1.50 min ?
Answer #1

B) Forward force on the truck = 4500 N

Backward force on the truck = 760 N

Net force on the truck = (4500-760) = 3740 N

By Newton's 2nd law, Fnet = mT.aT

Or, 3740 = 2200*aT

So, truck's acceleration, aT = (3740/2200) m/s^2

= 1.7 m/s^2